OpenStudy (anonymous):

3c + 2 = 4c - 9. 3c + 2 = 4c - 9. @Mathematics

OpenStudy (hoblos):

c=11

OpenStudy (anonymous):

combine like terms. 2+9=4c-3c

OpenStudy (anonymous):

0.25(7 + 3m) = -0.125 m im confused

OpenStudy (anonymous):

trust me,i want the help

OpenStudy (hoblos):

m=-2

OpenStudy (anonymous):

first divide the -0.125 by 0.25 that will leave you 7+3m==-0.5m

OpenStudy (anonymous):

then collect like terms. 7=-3.5m

OpenStudy (anonymous):

i would i do problems like A/4-9=5

OpenStudy (anonymous):

first move the 9, then you can take care of the 4

OpenStudy (anonymous):

like switch the 2 of them??

OpenStudy (anonymous):

no. move the -9 to the other side f the =. a/4=5+9

OpenStudy (anonymous):

its called collecting like terms

OpenStudy (anonymous):

i would i solve A/4 then?

OpenStudy (anonymous):

to move the 4 multiply both sides by 4. what you do to one side you do to the other.